Financial statements reflect only book values of the data that analysts use to evaluate a companys performance. To determineYou have taken a job as an entry-level analyst, and your boss has asked you to find the expected value of Adams Furnitures s763.00 - $20.25 $70.88 -$27.00 -$22.50 78.75 - $24.75 be the chat
Answer #1

1. (All in dollars)

Net sales 80000000
(less)Total operating costs 52000000
Operating income (EBIT) 28000000
(less)Tax 11200000
NOPAT 16800000

NOPAT (Net operating profit after taxes) = Operating income * (1-tax rate) = $28,000,000*(1-40%) = $16,800,000

EVA(Economic value added) = NOPAT - Invested capital*Cost of capital (after tax) = 16,800,000 - 83,500,000*10% = $8,450,000 (2nd option)

2. Expected value of stock = D1/(r-g) r = 10% (Return on equity), g = 5% (growth)

D1 is dividends expected to be paid in next year = Do*(1+growth%) = $3.75*(1+5%) = $3.75*1.05 = $3.9375

Expected value of stock = $3.9375/(10%-5%) = $78.75 (4th option)

3. If r = 12% (instead of 10% earlier)

Expected value of stock = $3.9375/(12%-5%) = $56.25

So Net change = $56.25 - $78.75 = - $22.5 (a decrease of $22.5)